Displacement of a particle at Time ’t’ = b+ b1t + b2t. Find the acceleration. I dont know how to solve this.. Please help me... Please!!

Grade:11

1 Answers

Vikas TU
7 years ago
Remember,
velocity is the differentiation of displacemtn with time 
accln. is the differentiation of velocity with time.
Thus,
v = dx/dt  and a = dv/dt.
thus,
v = b1 +2b2t
and 
a = 2b2 m/s^2